# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ IV-A4 Public Safety
Natural disaster prevention utilizes GI to aid people in mitigating the impact of unexpected extreme hazards, including earthquakes, wildfires, floods, and hurricanes. Bilal et al. [33] presents a batch normalization GCN network, aimed at early earthquake detection. It integrates a CNN with a GNN, where the CNN processes multi-station and three-component waveform data, and the extracted features are combined with location information for input into the GNN. The model employs batch normalization to reduce training epochs and stabilize activation value distributions, enhancing training efficiency and prediction accuracy for key earthquake parameters like magnitude, depth, and location. This approach of integrating CNNs with GCNs for efficient feature extraction and spatial relationship modeling is not limited to seismology. Similarly, in a related application, Jin et al. [138] tackle the challenge of predicting urban fire dynamics, which is crucial for urban safety and emergency response. They propose an innovative model that combines CNNs with GCNs, where CNNs extract pixel-level latent representations from fire situation awareness images and GCNs manage the spatial relationships between different urban areas. By integrating the strengths of CNNs and GCNs, Jin et al. effectively illustrate how this combined approach can address complex spatio-temporal phenomena, highlighting its versatility and potential across different domains.
@@
Urban crime is another critical factor imperiling residents’ properties and lives. To raise the attention of people potentially in criminal danger, researchers develop GI models to predict crime frequency in dedicated regions. Xia et al. [139] present the Spatial-Temporal Sequential Hypergraph network (ST-SHN), aimed at improving crime prediction by addressing the dynamic nature of criminal patterns in both spatial and temporal domains, and the evolving dependencies between different crime types. ST-SHN employs a graph-structured message-passing architecture combined with hypergraph learning to manage the spatial-temporal dynamics and global context of crime. In parallel, Sun et al. [140] present AGI-STAN, a framework that improves crime prediction by discarding reliance on domain-specific knowledge and predefined graphs. It uses adaptive graph learning to autonomously discern interdependencies among urban communities and integrates a time-aware self-attention mechanism to model temporally varying crime incidents. Both ST-SHN and AGI-STAN utilize GI to understand crime patterns, emphasizing the importance of spatial and temporal factors. ST-SHN focuses on global context through hypergraph learning, while AGI-STAN discovers community interdependencies and processes temporal information adaptively.
@@
Epidemic outbreaks, notably the recent novel coronavirus, pose significant threats to global public health systems. Accurate epidemic prediction is crucial, as it plays a vital role in safeguarding public health by informing and guiding proactive strategies to mitigate the impact of such health crises. In this context, GI have emerged as a powerful tool for modeling the complex interactions inherent in epidemic data. Both Kapoor et al. [141] and Keicher et al. [142] leverage the representational power of GNNs to address different facets of COVID-19 prediction, underscoring the versatility and efficacy of graph-based modeling in this domain. While both studies employ GNNs to harness the complexity of COVID-19 data, they differ in their specific applications and data integration strategies. Kapoor et al. focus on macro-level predictions of case numbers across regions, emphasizing the epidemiological aspect of the disease spread. In contrast, Keicher et al. concentrate on micro-level patient outcomes within a clinical setting, emphasizing the individual variability in disease progression and response to treatment.
IV-B Intelligent Robots and Vehicles
@@
IV-B1 Body Control and Manipulation
@@
With a model of the robot, we may envisage to command it to firmly grasp or manipulate some objects. This requires, beyond body modeling, to characterize the interaction between robots, objects, and environments. In this context, Li et al. [145] make a significant contribution by introducing dynamic particle interaction networks (DPINets), a differentiable, particle-based simulator. This innovative approach leverages dynamic interaction graphs to adeptly handle the manipulation of a variety of materials, including fluids, deformables, and rigid bodies. Building upon the foundation laid by the DPINets’ focus on material interaction, Xie et al. [146] present a deep imitation learning framework for bimanual robotic manipulation in continuous state-action spaces, focusing on the challenge of generalizing manipulation skills to objects in varying locations. Further extending the application of graph-based learning in robotics, Liang et al. [147] introduce a self-supervised learning method, target-oriented Deep Q-Network (DQN), which employs visual affordance graphs for a complex object grasping task, guiding robot actions using environmental cues in both simulated and real-world scenarios. These three works collectively exemplify the power of graph-based modeling in advancing robotic manipulation capabilities. From the dynamic interaction graphs of DPINets for simulating material interactions, to the hierarchical decomposition and RGNNs in the deep imitation learning framework, and finally to the self-supervised learning using visual affordance graphs.
IV-B2 Motion Prediction and Planning
@@
In many circumstances, robots are not launched individually but collaboratively, which requires global planning of robots to attain mutual goals [148]. For these cases, GI is also profitable with robots modeled in graphs. For instance, Fig. 7(d) shows a drone swarm committed to collision-freely flying from a starting point to a targeted location, where they can be viewed as a graph with links indicating their communication paths. Li et al. [149] tackle the challenge of decentralized multi-robot path planning by introducing a model that synthesizes local communication and decision-making policies for robots in constrained workspaces. The model combines a CNN for extracting features from local observations and a GNN for efficient communication among robots. The work is later extended in [150], which introduces the message-aware graph attention network to allow for more effective prioritization of important information from neighboring robots, showcasing improved performance and adaptability in large-scale autonomous mobile robot systems. A similar structure is exploited by [151], which presents the dynamic motion planning model based on graph neural networks and historical information to enhance path-finding in decentralized multi-robot systems.
IV-B3 Environment Exploration
@@
Contrary to motion planning where destinations are known, multi-robot exploration tasks demand exploring an unknown environment a key aspect of automation applications like cleaning, searching, and rescue. For these tasks, scheduling robot swarms to efficiently cover the underexplored field while avoiding conflicts is vital [152, 153]. To achieve that, researchers define the spatial relationships between robots and regions of the environment and intend to apply GI to regulate this complex process. Luo et al. [154] were among the pioneers in this domain, introducing a graph-based DRL approach for autonomous exploration in unknown territories. By translating the exploration task into a graph domain through a hierarchical map segmentation, they paved the way for utilizing a GCN to effectively assign exploration targets to agents. This seminal work highlighted the potential of GNNs in enhancing multi-agent exploration efficiency. Building upon this, Gosrich et al. [155] advanced the application of GNNs by developing a decentralized control policy for multi-robot sensor coverage. Their innovative approach allowed robots with limited sensing capabilities to efficiently detect events in regions of varying importance. In parallel, Zhang et al. [156] have taken the application of GNNs a step further with the introduction of Hierarchical-Hops Graph Neural Networks (H2GNN) for multi-robot coarse-to-fine exploration. Their approach stands out by enabling selective integration of key environmental information through a multi-head attention mechanism, which discerns the importance of information from different hops. Together, these studies provide a comprehensive view of how GNNs can be leveraged to address a spectrum of challenges in multi-agent robotics, setting the stage for future advancements in the field.
IV-C Human Sensing and Analysis
@@
Human sensing are core application in many edge services, which use edge devices to collect and process data related to human activities and behaviors. Various platforms are involved, including sensors, smartphones, wearables, cameras, etc., to monitor various aspects of human behavior such as movement, gestures, vital signs, facial expressions, and environmental interactions. GI models are employed in this context for those data that can be formulated in graphs. In general, these human sensing applications are mainly recognized in three types: affective computing with facial graphs, action recognition with human skeletons, and smart health with human-centric sensor networks.
IV-C1 Affective Computing
@@
Facial expressions contribute to more than 55% of the information perceived by individuals during verbal communication [157, 158], which plays a crucial role in conveying significant information related to emotional states and reactions within human interactions. To enable user-centric services deployed at the edge, the community has developed various intelligent affective computing methods to understand users’ facial affects. Notably, some researchers establish graph representations for faces, where facial landmarks are marked as vertices and are connected in facial shapes, and introduce GI models to reason users’ states. In a harmonious convergence, Xie et al. [159], Liu et al. [160], and Zhao et al. [161] have each harnessed the expressive power of GNNs to decode the complex language of human emotions manifested through facial movements. Specifically, Xie et al. [159] delve into the micro-level of facial expressions, presenting an innovative approach that harnesses the power of graph-based modeling to distinguish the subtleties of micro-expressions. They integrate action units with emotion category labels within a graph framework, thereby enhancing the precision of recognition tasks that are critical in various real-world applications. Building upon this graph-centric perspective, Liu et al. [160] propose a semantic graph-based dual-stream network for facial expression recognition (FER). This work addresses a key limitation of conventional CNN methods by incorporating semantic information into the recognition process. Their model not only learns from the physical appearance but also understands the deeper meaning behind facial movements, providing a more comprehensive analysis of emotional states. Similarly, Zhao et al. [161] contribute to the field by developing a geometry-aware FER framework. Their approach cleverly combines the structural analysis capabilities of a GNN with the feature extraction prowess of a CNN. This dual-pronged strategy allows for a detailed examination of both the geometric and appearance-based aspects of facial expressions, leading to a robust recognition system that is sensitive to the minute details of human emotions.
IV-C2 Action Recognition
@@
The rapid advancement of Human Action Recognition (HAR) and tracking techniques has emerged as a crucial catalyst for diverse edge applications, e.g., for security surveillance and robotic imitation. The rationale behind using GI models for action recognition is to model human bodies into skeleton-based graphs, as exemplified in Fig. 7(e). Specifically, body joints are regarded as vertices with each attached a feature vector indicating its 3D/2D coordinates and confidence scores, while they are connected along with the human skeletons. In the realm of graph-based models for skeleton dynamics, a spectrum of approaches has been explored, categorized by their focus on spatial or spatio-temporal characteristics. The work of Wang et al. [162] exemplifies the spatial approach with their introduction of Graph-PCNN, a framework that refines keypoint localization through a two-stage process, enhancing the accuracy of human pose estimation. This model-agnostic framework underscores the importance of accurate initial localization, which is further improved through a graph pose refinement module. Building upon the spatial foundation, Peng et al. [163] take a leap into the spatio-temporal domain by employing NAS to evolve an adaptive GCN for HAR. Their approach cleverly integrates dynamic graph modules and multiple-hop connections, which not only bolsters the network’s capacity to capture the nuances of motion in skeleton data but also aligns with the spatial-temporal emphasis of Graph-PCNN, albeit in a different context. Similarly, Jin et al. [164] address the complexity of multi-person pose estimation by proposing a hierarchical graph grouping method. This work, like Graph-PCNN, leverages the power of graph structures but diverges by reformulating the human part grouping into a graph clustering task, offering a fresh perspective on the integration of graph theory in pose estimation.
IV-C3 Smart Health
@@
Health condition analysis integrates GI with sensing techniques to enable healthcare monitoring and intervention. The synergy of these technologies is exemplified in Dong et al.’s work [165], where they introduce an end-to-end GNN model. This model harnesses mobile sensing data to detect early signs of influenza-like symptoms by encapsulating the dynamics of state transitions and internal dependencies within human behaviors through graph representation. Building upon this foundation, the concept of semi-supervised learning is elegantly integrated in the work by Dong et al. [166], where they propose a Graph Instance Transformer (GIT). This novel model not only combines multiple instance learning with contrastive self-supervised learning but also demonstrates the adaptability of GNNs in predicting early signs of mental health disorders. GIT’s design is particularly noteworthy for its ability to make use of the extensive unannotated data across various levels of granularity, thereby enhancing the predictive power of machine learning models with limited supervision. Similarly, Jia et al. [167] contribute to the field by introducing GraphSleepNet, a GNN-based framework specifically tailored for automatic sleep stage classification. This framework adeptly tackles the challenge of utilizing brain spatial features and the transition information among sleep stages, showcasing the versatility of GNNs in addressing complex health-related problems.
IV-D Location-Based Services
@@ IV-D1 Social POI Recommendation
Social POI recommendation utilizes users’ activities on social media platforms to provide recommendations for friends they may know or products they may like to purchase. In a pioneering approach, Kefalas et al. [168] enhance the traditional recommender system by introducing a hybrid tripartite graph that intertwines user, location, and session networks. This structure is adeptly analyzed using an advanced random walk with restart algorithm, which adeptly captures the fluid nature of user preferences over time and space. Building upon this foundation, Salamat et al. [169] introduce HeteroGraphRec, a sophisticated recommender system that further sophisticates the social network model by considering it as a heterogeneous graph. Continuing this thread of innovation, Wang et al. [170] tackle the specific challenge of session-based recommendations by proposing a spatio-temporal GNN model. This model adeptly predicts a user’s immediate interests by considering the sequence of their recent anonymous behaviors, thereby providing a more nuanced and responsive recommendation strategy.
IV-D2 Geographical POI Recommendation
@@
The objective of geographical POI recommendation is to exploit users’ behaviors to predict their potentially interested locations. Both Luo et al. [171] and Yang et al. [172] have introduced models that ingeniously incorporate spatio-temporal dynamics to predict user preferences for future locations. A common thread between these works is their reliance on attention mechanisms to dissect the complex interplay of spatial and temporal factors. Luo et al.’s STAN model introduces self-attention to directly model the spatial and temporal aspects of user check-ins, offering detailed insights into user behavior patterns. This approach excels at capturing the significance of non-adjacent and non-consecutive interactions. In a parallel yet distinctive approach, Yang et al.’s STAM model enhances GNNs by incorporating scaled dot-product and multi-head attention mechanisms. STAM enriches the learning of neighbor embeddings, providing a broader perspective on user-item interactions. The transition from STAN to STAM illustrates the field’s maturation, moving from specific spatio-temporal interactions to a more integrated and comprehensive understanding of user preferences in recommendation systems.
@@
Figure 8: Example point cloud processing with GI models. The point cloud data captured by autonomous vehicles can be reorganized in a graph format and processed with GI models for high-precision object detection. Image adopted from [173].
IV-E Graph-Based Mobile Vision

### v2

@@ V-A4 Public Safety
Natural disaster prevention utilizes GI to aid people in mitigating the impact of unexpected extreme hazards, including earthquakes, wildfires, floods, and hurricanes. Bilal et al. [28] presents a batch normalization GCN for early earthquake detection, which integrates a CNN with a GNN. The model employs batch normalization to reduce training epochs and stabilize activation value distributions, enhancing training efficiency and prediction accuracy for key earthquake parameters like magnitude, depth, and location. This approach of integrating CNNs with GCNs for efficient feature extraction and spatial relationship modeling is not limited to seismology. Similarly, Jin et al. [128] tackle the challenge of predicting urban fire dynamics, which is crucial for urban safety and emergency response. They propose a model that combines CNNs with GCNs, where CNNs extract pixel-level latent representations from fire situation awareness images and GCNs manage the spatial relationships between different urban areas. By integrating the strengths of CNNs and GCNs, [128] effectively illustrates how this combined approach can address complex spatio-temporal phenomena, highlighting its versatility and potential across different domains.
@@
Urban crime is another critical factor imperiling residents’ properties and lives. To raise the attention of people potentially in criminal danger, researchers develop GI models to predict crime frequency in dedicated regions. Xia et al. [129] present the Spatial-Temporal Sequential Hypergraph network (ST-SHN), aimed at improving crime prediction by addressing the dynamic nature of criminal patterns in both spatial and temporal domains, and the evolving dependencies between different crime types. ST-SHN employs a graph-structured message-passing architecture combined with hypergraph learning to manage the spatial-temporal dynamics and global context of crime. In parallel, Sun et al. [130] present AGI-STAN, a framework that improves crime prediction by discarding reliance on domain-specific knowledge and predefined graphs. It uses adaptive graph learning to autonomously discern interdependencies among urban communities and integrates a time-aware self-attention mechanism to model temporally varying crime incidents. Toward the same problem, these two methods take different optimization paths: ST-SHN focuses on global context through hypergraph learning, while AGI-STAN discovers community interdependencies and processes temporal information adaptively.
@@
Epidemic outbreaks, notably the recent novel coronavirus, pose significant threats to global public health systems. Accurate epidemic prediction is crucial, as it plays a vital role in safeguarding public health by informing and guiding proactive strategies to mitigate the impact of such health crises. In this context, GI has emerged as a powerful tool for modeling the complex interactions inherent in epidemic data. Both Kapoor et al. [131] and Keicher et al. [132] leverage the expressiveness power of GNNs to address different facets of COVID-19 prediction, underscoring the versatility and efficacy of graph-based modeling in this domain. While both studies employ GNNs to harness the complexity of COVID-19 data, they differ in their specific applications and data integration strategies. Kapoor et al. focus on macro-level predictions of case numbers across regions, emphasizing the epidemiological aspect of the disease spread. In contrast, Keicher et al. concentrate on micro-level patient outcomes within a clinical setting, emphasizing the individual variability in disease progression and response to treatment.
V-B Intelligent Robots and Vehicles
@@
V-B1 Body Control and Manipulation
@@
With a model of the robot, we may envisage to command it to firmly grasp or manipulate some objects. This requires, beyond body modeling, to characterize the interaction between robots, objects, and environments. In this context, Li et al. [135] make a significant contribution by introducing dynamic particle interaction networks (DPINets), a differentiable, particle-based simulator. This innovative approach leverages dynamic interaction graphs to adeptly handle the manipulation of a variety of materials, including fluids, deformables, and rigid bodies. Building upon the foundation laid by the DPINets’ focus on material interaction, Xie et al. [136] present a deep imitation learning framework for bimanual robotic manipulation in continuous state-action spaces, focusing on the challenge of generalizing manipulation skills to objects in varying locations. Further extending the application of graph-based learning in robotics, Liang et al. [137] introduce a self-supervised learning method, target-oriented Deep Q-Network (DQN), which employs visual affordance graphs for a complex object grasping task, guiding robot actions using environmental cues in both simulated and real-world scenarios. These three works collectively exemplify the power of graph-based modeling in advancing robotic manipulation capabilities
V-B2 Motion Prediction and Planning
@@
In many circumstances, vehicles are not launched individually but collaboratively, which requires global planning of vehicles to attain mutual goals [138]. For these cases, GI is also profitable with vehicles modeled in graphs. For instance, Fig. 7(d) shows a drone swarm committed to collision-freely flying from a starting point to a targeted location, where they can be viewed as a graph with links indicating their communication paths. Another illustration is shown in Fig. 8, where the movement of vehicles is modeled in a time-varying graph and fed into a GAE for learning graph representations. Through a predictor and a planner, the EGI system can export the prediction and planning of the vehicle’s next motion. Li et al. [139] tackle the challenge of decentralized multi-robot path planning by introducing a model that synthesizes local communication and decision-making policies for robots in constrained workspaces. The model combines a CNN for extracting features from local observations and a GNN for efficient communication among robots. A similar structure is exploited by [140], which presents the dynamic motion planning model based on graph neural networks and historical information to enhance path-finding in decentralized multi-robot systems.
Figure 8: An illustration of vehicle trajectory planning via GI models, where the historical trajectories of vehicles are modeled as a time-varying graph.
V-B3 Environment Exploration
Contrary to motion planning where destinations are known, multi-robot exploration tasks demand exploring an unknown environment a key aspect of automation applications like cleaning, searching, and rescue. For these tasks, scheduling robot swarms to efficiently cover the underexplored field while avoiding conflicts is vital [141, 142].
@@ V-B3 Environment Exploration
To achieve that, researchers define the spatial relationships between robots and regions of the environment and intend to apply GI to regulate this complex process. Luo et al. [143] were among the pioneers in this domain, introducing a graph-based DRL approach for autonomous exploration in unknown territories. By translating the exploration task into a graph domain through a hierarchical map segmentation, they paved the way for utilizing a GCN to effectively assign exploration targets to agents. This seminal work highlighted the potential of GNNs in enhancing multi-agent exploration efficiency. Building upon this, Gosrich et al. [144] advanced the application of GNNs by developing a decentralized control policy for multi-robot sensor coverage. Their innovative approach allowed robots with limited sensing capabilities to efficiently detect events in regions of varying importance. In parallel, Zhang et al. [145] have taken the application of GNNs a step further with the introduction of Hierarchical-Hops Graph Neural Networks (H2GNN) for multi-robot coarse-to-fine exploration. Their approach stands out by enabling selective integration of key environmental information through a multi-head attention mechanism, which discerns the importance of information from different hops.
Together, these studies provide a comprehensive view of how GNNs can be leveraged to address a spectrum of challenges in multi-agent robotics, setting the stage for future advancements in the field.
V-C Human Sensing and Analysis
@@
Human sensing is one of the core applications in many edge services, which use edge devices to collect and process data related to human activities and behaviors. Various platforms are involved, including sensors, smartphones, wearables, cameras, etc., to monitor various aspects of human behavior such as movement, gestures, vital signs, facial expressions, and environmental interactions. GI models are employed in this context for those data that can be formulated in graphs. In general, these human sensing applications are mainly recognized in three types: affective computing with facial graphs, action recognition with human skeletons, and smart health with human-centric sensor networks.
V-C1 Affective Computing
Facial expressions contribute to more than 55% of the information perceived by individuals during verbal communication [146, 147], which plays a crucial role in conveying significant information related to emotional states and reactions within human interactions. To enable user-centric services deployed at the edge, the community has developed various intelligent affective computing methods to understand users’ facial affects. Notably, some researchers establish graph representations for faces, where facial landmarks are marked as vertices and are connected in facial shapes, and introduce GI models to reason users’ states.
@@ V-C1 Affective Computing
In a harmonious convergence, Xie et al. [148], Liu et al. [149], and Zhao et al. [150] have each harnessed GNNs to decode the complex language of human emotions manifested through facial movements. Specifically, Xie et al. [148] delve into the micro-level of facial expressions, integrating action units with emotion category labels within a graph framework, thereby enhancing the precision of recognition tasks that are critical in various real-world applications. Building upon this graph-centric perspective, Liu et al. [149] propose a semantic graph-based dual-stream network that addresses a key limitation of conventional CNN methods by incorporating semantic information into the recognition process. Their model not only learns from physical appearance but also understands the deeper meaning behind facial movements. Similarly, Zhao et al. [150] contribute to the field by developing a geometry-aware FER framework, which combines the structural analysis capabilities of a GNN with the feature extraction prowess of a CNN. This dual-pronged strategy allows for a detailed examination of both the geometric and appearance-based aspects of facial expressions, leading to a robust recognition system that is sensitive to the minute details of human emotions.
V-C2 Action Recognition
The rapid advancement of Human Action Recognition (HAR) and tracking techniques has emerged as a crucial catalyst for diverse edge applications, e.g., for security surveillance and robot imitation. The rationale behind using GI models for action recognition is to model human bodies into skeleton-based graphs, as exemplified in Fig. 7(e). Specifically, body joints are regarded as vertices with each attached a feature vector indicating its 3D/2D coordinates and confidence scores, while they are connected along with the human skeletons.
@@ V-C2 Action Recognition
In the realm of graph-based models for skeleton dynamics, a spectrum of approaches has been explored, categorized by their focus on spatial or spatio-temporal characteristics. Wang et al. [151] present Graph-PCNN, a framework that refines keypoint localization through a two-stage process, enhancing the accuracy of human pose estimation. This model-agnostic framework underscores the importance of accurate initial localization, which is further improved through a graph pose refinement module. Building upon the spatial foundation, Peng et al. [152] take a leap into the spatio-temporal domain by employing NAS to evolve an adaptive GCN for HAR. Their approach integrates dynamic graph modules and multiple-hop connections, which not only bolsters the network’s capacity to capture the nuances of motion in skeleton data but also aligns with the spatial-temporal emphasis of Graph-PCNN, albeit in a different context. Similarly, Jin et al. [153] address the complexity of multi-person pose estimation by proposing a hierarchical graph grouping method. This work, like Graph-PCNN, leverages the power of graph structures but diverges by reformulating the human part grouping into a graph clustering task, offering a fresh perspective on the integration of graph theory in pose estimation.
V-C3 Smart Health
@@
Health condition analysis integrates GI with sensing techniques to enable healthcare monitoring and intervention. The synergy of these technologies is exemplified in Dong et al.’s work [154], where they introduce a GI model that harnesses mobile sensing data to detect early signs of influenza-like symptoms by encapsulating the dynamics of state transitions and internal dependencies within human behaviors through graph representation. Building upon this foundation, the concept of semi-supervised learning is elegantly integrated in the work [155], where they propose a Graph Instance Transformer (GIT). GIT not only combines multi-instance learning with contrastive self-supervised learning but also demonstrates the adaptability of GNNs in predicting early signs of mental health disorders. Similarly, Jia et al. [156] contribute to the field by introducing GraphSleepNet, a GI framework specifically tailored for automatic sleep stage classification. This framework adeptly tackles the challenge of utilizing brain spatial features and the transition information among sleep stages, showcasing the versatility of GNNs in addressing complex health-related problems.
V-D Location-Based Services
@@ V-D1 Social POI Recommendation
Social POI recommendation utilizes users’ activities on social media platforms to provide recommendations for friends they may know or products they may like to purchase. In a pioneering approach, Kefalas et al. [157] enhance the traditional recommender system by introducing a hybrid tripartite graph that intertwines user, location, and session networks. This structure is analyzed using an advanced random walk with a restart algorithm, which captures the fluid nature of user preferences over time and space. Building upon this foundation, Salamat et al. [158] introduce HeteroGraphRec, a sophisticated recommender system that considers social networks a heterogeneous graph. Continuing this thread of innovation, Wang et al. [159] tackle the specific challenge of session-based recommendations by proposing an STGCN model. This model predicts a user’s immediate interests by considering the sequence of their recent anonymous behaviors, thereby providing a more nuanced and responsive recommendation strategy.
V-D2 Geographical POI Recommendation
@@
The objective of geographical POI recommendation is to exploit users’ behaviors to predict their potentially interested locations. Both Luo et al. [160] and Yang et al. [161] have introduced models that ingeniously incorporate spatio-temporal dynamics to predict user preferences for future locations. A common thread between these works is their reliance on attention mechanisms to dissect the complex interplay of spatial and temporal factors. Luo et al.’s STAN model [160] introduces self-attention to directly model the spatial and temporal aspects of user check-ins, excelling at capturing the significance of non-adjacent and non-consecutive interactions. In a parallel yet distinctive approach, Yang et al.’s STAM model enhances GNNs by incorporating scaled dot-product and multi-head attention mechanisms, which enriches the learning of neighbor embeddings. The transition from STAN to STAM illustrates the field’s maturation, moving from specific spatio-temporal interactions to a more integrated and comprehensive understanding of user preferences in recommendation systems.
@@
Figure 9: Example point cloud processing with GI models. The point cloud data captured by autonomous vehicles can be reorganized in a graph format and processed with GI models for high-precision object detection. Image adapted from [162].
V-E Graph-Based Mobile Vision